OpenStudy (anonymous):

Add or subtract the rational expressions, as indicated, and simplify the answer.

OpenStudy (anonymous):

\[\frac{x^2-1}{x-2}-\frac{x-1}{x+1}\]

OpenStudy (anonymous):

your only choice for the denominator is the product of these two so it will be \[(x-2)(x+1)\] and your work will be in the numerator start with \[\frac{(x^2-1)(x+1)-(x-1)(x-2)}{(x-2)(x+1)}\] then multiply out in the numerator, combine like terms, etc. do not look to cancel, it is very unlikely

OpenStudy (anonymous):

Oh i see well thanks. I have one more if you would like to help me solve it.

OpenStudy (anonymous):

sure when all is said and done your numerator should be \[x^3+2 x-3\] or if you want to factor \[(x-1) (x^2+x+3)\] so you will get \[\frac{x^3+2x-3}{(x-2)(x+1)}\]

OpenStudy (anonymous):

So the final answer is the bottom one?

OpenStudy (anonymous):

yes, unless you want to write the numerator in fatored form

OpenStudy (anonymous):

no it just says to write it in simplified form.

OpenStudy (anonymous):

One sec i will write the other problem.

OpenStudy (anonymous):

\[\frac{y+2}{y-2}+\frac{y-6}{y+4}+\frac{4}{y^2+2y-8}\]

OpenStudy (anonymous):

if you are lucky \[y^2+2y-8\] is the product of \[(x-2)(x+4)\] and it is, so you can use that one as your least common denominator

OpenStudy (anonymous):

\[\frac{y+2}{y-2}+\frac{y-6}{y+4}+\frac{4}{y^2+2y-8}\] \[\frac{y+2}{y-2}\times \frac{y+4}{y+4}+\frac{y-6}{y+4}\times \frac{y-2}{y-2}+\frac{4}{y^2+2y-8}\]

OpenStudy (anonymous):

now your denominator is the same for each fraction, namely \[(y+4)(y-2)\] or \[y^2+2y-8\] and again all the work is in the numerator. you have to multiply out, combine like terms etc.

OpenStudy (anonymous):

\[(y+2)(y+4)+(y-6)(y-2)+4\] is your numerator. when you do the algebra you should get \[2 y^2-2 y+24\] for the numerator

OpenStudy (anonymous):

Thanks man your a genius lol.

OpenStudy (anonymous):

Just making sure the final answer is the bottom one?

OpenStudy (anonymous):

your only choice for the denominator is the product of these two so it will be (x−2)(x+1) and your work will be in the numerator start with (x2−1)(x+1)−(x−1)(x−2)(x−2)(x+1) then multiply out in the numerator, combine like terms, etc. do not look to cancel, it is very unlikely 22 minutes agoReport Abuse 11 alexsu22 This is the asker Oh i see well thanks. I have one more if you would like to help me solve it. 21 minutes agoReport Abuse 100 satellite73 1 Good Answer sure when all is said and done your numerator should be x3+2x−3 or if you want to factor (x−1)(x2+x+3) so you will get x3+2x−3(x−2)(x+1) 20 minutes agoReport Abuse 11 alexsu22 This is the asker So the final answer is the bottom one? 19 minutes agoReport Abuse 100 satellite73 1 Good Answer yes, unless you want to write the numerator in fatored form 18 minutes agoReport Abuse 11 alexsu22 This is the asker no it just says to write it in simplified form. 18 minutes agoReport Abuse 11 alexsu22 This is the asker One sec i will write the other problem. 16 minutes agoReport Abuse 11 alexsu22 This is the asker y+2y−2+y−6y+4+4y2+2y−8 12 minutes agoReport Abuse 100 satellite73 1 Good Answer if you are lucky y2+2y−8 is the product of (x−2)(x+4) and it is, so you can use that one as your least common denominator 11 minutes agoReport Abuse 100 satellite73 1 Good Answer y+2y−2+y−6y+4+4y2+2y−8 y+2y−2×y+4y+4+y−6y+4×y−2y−2+4y2+2y−8 10 minutes agoReport Abuse 100 satellite73 1 Good Answer now your denominator is the same for each fraction, namely (y+4)(y−2) or y2+2y−8 and again all the work is in the numerator. you have to multiply out, combine like terms etc. 9 minutes agoReport Abuse 100 satellite73 1 Good Answer (y+2)(y+4)+(y−6)(y−2)+4 is your numerator. when you do the algebra you should get 2y2−2y+24 for the numerator 8 minutes agoReport Abuse 11 alexsu22 This is the asker Thanks man your a genius lol. 4 minutes agoReport Abuse 11 alexsu22 This is the asker Just making sure the final answer is the bottom one? 3 minutes agoReport Abuse Type your reply EquationDrawAttach File

OpenStudy (anonymous):

i know im a genius alexsu :) U R TOO NO WORRIES K

OpenStudy (anonymous):

i am genius

OpenStudy (anonymous):

Everybody is genius except me :/